OpenStudy (anonymous):

s=t^3-5t^2+8, s is the displacement of an object, find an expression for the velocity v m/s, of the particle after 't' seconds. and also find the time at which the acceleration of the particle is 20m/s^2 ......

OpenStudy (anonymous):

\[s(t)=t^3-5t^2+8\] \[v(t)=s'(t)=3t^2-10t\]

OpenStudy (anonymous):

\[a(t)=v'(t)=6t-10\] set \[a(t)=6t-10=20\] \[6t=30\] \[t=5\]